# Equity auction dynamics: latent liquidity models with activity acceleration

### v1

@@ I.1 Model description
Our starting point is the coupled reaction-diffusion equations for the revealed $\rho^{(r)}$ and the latent $\rho^{(l)}$ orders densities derived in Dall’Amico et al. [20]. By allowing for general price and time dependencies, the buy side equations read
@@
We complement Eqs (1) and (2) with a set of boundary conditions. For the sell side, we assume that the slope of the latent order book has a non-zero limit for large $x$, $\partial_{x}\rho_{S}^{(l)}\underset{x\rightarrow+\infty}{\longrightarrow}a>0$, and that $\rho_{S}^{(l)}$ does not diverge when $x\rightarrow-\infty$. We also impose that the revealed order book $\rho_{S}^{(l)}$ does not diverge for large prices $|x|\rightarrow+\infty$. Similar boundary conditions hold for the buy side.
Dall’Amico et al. [20] posit the existence of interactions between the latent order book $\rho^{(l)}$ and the revealed order book $\rho^{(r)}$. The reveal rate $\nu_{r}\Gamma_{r}$ is the rate at which latent trading intentions in the latent order book materialize into actual orders in the revealed order book and the unreveal rate $\nu_{l}\Gamma_{l}$ is the rate at which actual orders revert back to latent intentions. Assuming the absence of exogenous depositions and cancellations in the latent order book, an order submission is equivalent to an order revelation, and similarly, an order cancellation is equivalent to an order unrevelation.
@@
Lastly, in open markets, the current price $p_{t}$ is the point where vanishing supply meets vanishing demand; accordingly, Dall’Amico et al. [20] define $p_{t}$ as the point where $p\rightarrow\left(\rho_{B}^{(r)}-\rho_{S}^{(r)}\right)(p)$ changes sign. In auctions instead, the indicative price $p_{t}$ is determined by equalizing supply and demand $S(p_{t})=D(p_{t})$ where $S(p)=\int_{p<p^{\prime}}\rho_{S}^{(r)}(p^{\prime})\mathrm{d}p^{\prime}$ and $D(p)=\int_{p>p^{\prime}}\rho_{B}^{(r)}(p^{\prime})\mathrm{d}p^{\prime}$. When centering equations around $\log(p_{t})$, we implicitly assume that the indicative price evolves independently of order densities. However, changes in the sell (resp. buy) order density for $x<0$ (resp. $x>0$) directly influence supply (resp. demand) subsequently changing the indicative price. We disregard this effect and consider that the dynamics of the indicative price are independent of average order densities around the indicative price. In the following, we focus on the sell-side equations (2), and denote $\rho_{S}$ as $\rho$ when there is no ambiguity.
@@ I.2 A stationary solution
We solve Eq. (4) separately on $\mathbb{R}^{+}$ and $\mathbb{R}^{-}$ as the first derivative of $x\rightarrow\Gamma_{r/l}(x)$ might not be continuous at $x=0$. Thus, whenever $D_{l}\neq 0$, the latent order book should be linear. Incorporating the boundary conditions $\partial_{x}\rho^{(l)}\underset{x\rightarrow+\infty}{\longrightarrow}a>0$ and $\partial_{x}\rho^{(l)}\underset{x\rightarrow-\infty}{\longrightarrow}b\geq 0$ yields
@@ II.2.3 Measuring diffusion in the revealed order book: $D_{r}$
where $\sigma$ is the volatility of the indicative price, $\text{Var}(\beta)$ is a prefactor that encodes the heterogeneity of agent reactions to price movements, and $\Gamma_{D}(x,y)$ is the rate at which orders are updated from log-price $x$ to log-price $y$. Should agents neither over-react nor under-react to indicative price movements ($\beta=1$ for all agents), the first term vanishes. Tóth et al. [44] consider a unit prefactor $\text{Var}(\beta)=1$. The definition of Eq. (24) corresponds to that of the revealed diffusion coefficient $D_{r}$ in our framework, provided we only account for visible price reassessments. Note that the under-diffusive nature of the indicative price may result in a different first-term contribution to the diffusion coefficient.
@@ II.3 Solving the full dynamical equations
$(\nu_{r}\Gamma_{r})(x,t)=\begin{cases}\frac{w\cdot C_{r}}{\gamma_{r}+T-\max\{t,t_{r}^{(0)}\}}\cdot e^{-x/x_{r}}+\frac{(1-w)\cdot C_{r}}{\gamma_{r}+T-t_{r}^{(0)}}\cdot e^{-x/(k\cdot x_{r})}\quad&,x\geq 0;\\ A^{*}\cdot e^{-x/x_{r}^{*}}&,-x_{0}\leq x<0;\\ m\cdot\frac{C_{r}}{\gamma_{r}}&,x<-x_{0},\end{cases}$ (26)
@@ III The anomalous scaling of the indicative price
Chen et al. [19] show that $H$ can be decomposed as $H=J+L+M+1$, where each of the exponents $J,L,M$ is associated with the failure of one condition of the central limit theorem: (i) the presence of correlations/long-term memory, (ii) infinite variance, and (iii) the non-stationarity of increments.
@@
  • the increments of the indicative price do not exhibit long-term memory during the accumulation period as $J\approx 1/2$ for most stocks and regimes. Therefore, the indicative price is efficient in the sense of Chen et al. [19]. Note that as the auction end approaches, the rescaled range statistic $R_{t}/S_{t}$ flattens due to the stabilization of the indicative price. As a result, $J$ is not defined in the last regime;

### v2

@@ I.1 Model description
Our starting point is the coupled reaction-diffusion equations for the revealed $\rho^{(r)}$ and the latent $\rho^{(l)}$ orders densities derived in Dall’Amico et al. [20]: they posit the existence of interactions between the latent order book $\rho^{(l)}$ and the revealed order book $\rho^{(r)}$. The reveal rate $\nu_{r}\Gamma_{r}$ is the rate at which latent trading intentions in the latent order book materialize into actual orders in the revealed order book and the unreveal rate $\nu_{l}\Gamma_{l}$ is the rate at which actual orders revert back to latent intentions. In addition, a diffusion mechanism is included in both the revealed and the latent order book corresponding to price updates in each order book with diffusion coefficient $D_{r}$ and $D_{l}$, respectively.
Assuming the absence of exogenous depositions and cancellations in the latent order book, an order submission is equivalent to an order revelation, and similarly, an order cancellation is equivalent to an order unrevelation. By allowing for general price and time dependencies, the buy side equations read
@@
We complement Eqs (1) and (2) with a set of boundary conditions. Let us focus on the sell side for the time being. It makes sense to assume that the latent liquidity book $\rho_{S}^{(l)}$ is an increasing function of $x$, as more people are willing to sell assets at a higher price. This means that one can impose a boundary condition on the slope of the latent order book: $\partial_{x}\rho_{S}^{(l)}\underset{x\rightarrow+\infty}{\longrightarrow}a>0$, which corresponds to the latent liquidity parameter of Donier et al. [24], and that $\rho_{S}^{(l)}$ does not diverge when $x\rightarrow-\infty$, which reflects the fact that the number of people willing to sell at a vanishing price is not infinite. We also impose that the revealed order book $\rho_{S}^{(r)}$ does not diverge for large prices $|x|\rightarrow+\infty$. This means that agents tend not to reveal their reservation price when it is far away from the indicative price. In practice, agents can send market orders or matchable limit orders far away from the indicative price in order to guarantee their participation in the auction volume. Market orders are not included in the densities $\rho_{S}^{(r)}$ and $\rho_{B}^{(r)}$ as latent liquidity models are defined in the reference frame of the indicative price.
Similar boundary conditions hold for the buy side.
@@
Lastly, in open markets, the current price $p_{t}$ is the point where vanishing supply meets vanishing demand; accordingly, Dall’Amico et al. [20] define $p_{t}$ as the point where $p\rightarrow\left(\rho_{B}^{(r)}-\rho_{S}^{(r)}\right)(p)$ changes sign. In auctions instead, the indicative price $p_{t}$ is determined by equalizing supply and demand
$S(p_{t})+MO_{S,t}=D(p_{t})+MO_{B,t},$ (3)
where $S(p)=\int_{p^{\prime}<p}\rho_{S}^{(r)}(p^{\prime})\mathrm{d}p^{\prime}$, $D(p)=\int_{p^{\prime}>p}\rho_{B}^{(r)}(p^{\prime})\mathrm{d}p^{\prime}$, and $MO_{B}$ and $MO_{S}$ is the volume of buy and sell auction market orders. Equation (3) makes it clear that only the difference of market order volume has an influence on the indicative price. As latent liquidity models are defined in the reference frame of the indicative price, market orders are not included in the densities $\rho_{S}^{(r)}$ and $\rho_{B}^{(r)}$.
When centering equations around $\log(p_{t})$, we implicitly assume that the indicative price evolves independently of order densities. However, changes in the sell (resp. buy) order density for $x<0$ (resp. $x>0$) directly influence supply (resp. demand) subsequently changing the indicative price. We disregard this effect and consider that the dynamics of the indicative price are independent of average order densities around the indicative price. In the following, we focus on the sell-side equations (2), and denote $\rho_{S}$ as $\rho$ when there is no ambiguity.
@@ I.2 A stationary solution
We solve Eq. (5) separately on $\mathbb{R}^{+}$ and $\mathbb{R}^{-}$ as the first derivative of $x\rightarrow\Gamma_{r/l}(x)$ might not be continuous at $x=0$. Thus, whenever $D_{l}\neq 0$, the latent order book should be linear. Incorporating the boundary conditions $\partial_{x}\rho^{(l)}\underset{x\rightarrow+\infty}{\longrightarrow}a>0$ and $\rho^{(l)}\underset{x\rightarrow-\infty}{\longrightarrow}b\geq 0$ yields
@@ II.2.3 Measuring diffusion in the revealed order book: $D_{r}$
[…]
The definition of Eq. (25) corresponds to that of the revealed diffusion coefficient $D_{r}$ in our framework, provided we only account for visible price reassessments. Note that the under-diffusive nature of the indicative price may result in a different first-term contribution to the diffusion coefficient.
@@ II.3 Solving the full dynamical equations
$(\nu_{r}\Gamma_{r})(x,t)=\begin{cases}\frac{w\cdot C_{r}}{\gamma_{r}+T-\max\{t,t_{r}^{(0)}\}}\cdot e^{-x/x_{r}}+\frac{(1-w)\cdot C_{r}}{\gamma_{r}+T-t_{r}^{(0)}}\cdot e^{-x/(k\cdot x_{r})}\quad&,x\geq 0;\\ A^{*}\cdot e^{x/x_{r}^{*}}&,-x_{0}\leq x<0;\\ m\cdot\frac{C_{r}}{\gamma_{r}}&,x<-x_{0},\end{cases}$ (27)
@@ III The anomalous scaling of the indicative price
Chen et al. [19] show that $H$ can be decomposed as $H=J+L+M-1$, where each of the exponents $J,L,M$ is associated with the failure of one condition of the central limit theorem: (i) the presence of correlations/long-term memory, (ii) infinite variance, and (iii) the non-stationarity of increments.
@@
  • the increments of the indicative price do not exhibit long-term memory during the accumulation period as $J\approx 1/2$ for most stocks and regimes. Note that as the auction end approaches, the rescaled range statistic $R_{t}/S_{t}$ flattens due to the stabilization of the indicative price. As a result, $J$ is not defined in the last regime. We thus conclude that the indicative price is efficient in the sense of Chen et al. [19]: even if the apparent Hurst exponent is not $1/2$, $J=1/2$ precludes price predictions. Finally, [7] report an overreaction of the indicative price based on the following definition of the indicative jump on close $J(t)$
    $J(t)=\frac{p_{t}-p_{\text{ref}}}{p_{\text{auction}}-p_{\text{ref}}}.$ (37)
    However, this incorporates the first jump from $p_{\text{ref}}$ (the last price of the continuous trading phase) to $p_{0}$ (the first indicative price of the accumulation period). When the reference price is the first indicative price of the accumulation period $p_{0}$, $J(t)$ becomes $\tilde{J}(t)$
    $\tilde{J}(t)=\frac{p_{t}-p_{0}}{p_{\text{auction}}-p_{0}},$ (38)
    which does not display any overreaction. In addition, we find that the $p_{\text{auction}}-p_{\text{ref}}$ does not have systematically the same sign as $p_{\text{auction}}-p_{0}$ (only $\approx$50of the time). Thus, the first jump from $p_{\text{ref}}$ to $p_{0}$ is not predictive of the direction of the auction price.
@@ Acknowledgments
Disclosure of interest
The authors have no competing interests to declare.
Declaration of funding